OpenStudy (raden):

If f(x) = sin(sin(sin(...(sin(sin x))...))) then then value of f '(180°) = .....

OpenStudy (anonymous):

is that the derivative?

OpenStudy (raden):

f'(x) is the first derivative of f(x)

OpenStudy (anonymous):

i believe answer is 1

OpenStudy (raden):

how?

OpenStudy (anonymous):

i think its -1 and chain rule

OpenStudy (anonymous):

well after derivating all of ur outer functions are \(\cos\) lets try it with an example \[f(x)=\sin(\sin(\sin x)))\]\[f'(x)=[(\sin(\sin x))]' \ \cos(\sin(\sin x)))\]\[ \ \ \ \ \ \ \ = \cos x \ \cos(\sin x) \ \cos(\sin(\sin x))) \]\[f'(\pi)=\cos \pi \cos 0 \cos 0=-1\] so @completeidiot is right answer is -1............:)

OpenStudy (raden):

thanks mukushla what about with "chain rule", @completeidiot, i want study it...

OpenStudy (anonymous):

what i did was chain rule...

OpenStudy (raden):

hohoho... sorry, i just knowed

OpenStudy (anonymous):

np man....:)

OpenStudy (anonymous):

Haha this is confusing! You guys are smart! XD lol